Points A, B, and C are collinear. Point B is between A and C. Solve for X
AB=5, BC= 11+2x, AC=x+16
1
Expert's answer
2019-11-19T10:18:06-0500

Given, the value of AB =5

BC = 11 + 2x

AC = x + 16

Since, the points A, B and C are collinear

AC = AB + BC

Putting all the values, we get

x +16= 5 + (11+2x)

or, x +16= 5+ 11+2x

or, x+16=16+2x

or, 2x - x = 16-16

or, x = 0

Hence, the value of x is zero.